OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−6, 4) and (2, 0).

OpenStudy (anonymous):

There is more than one answer ...

OpenStudy (2bornot2b):

\[y-4=\frac{4-0}{-6-2}(x-(-6))\]

OpenStudy (anonymous):

slope = (4-0)/(-6 - 2) = 4/-8 = -1/2 Therefore y - 0 = (-1/2)(x - 2) or y - 4 = (-1/2)(x + 6)